OpenStudy (anonymous):

if y=3-u\2+u and u=4x\1-x2; find dy\dx.

OpenStudy (lgbasallote):

can you draw it?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

why not?

OpenStudy (anonymous):

okay first 2+u-3-u/(2+u)^2

OpenStudy (anonymous):

wrong

OpenStudy (lgbasallote):

oh..he knows the answer..fascinating

OpenStudy (anonymous):

answer can not match

OpenStudy (anonymous):

okay we use a differential

OpenStudy (anonymous):

it's right

OpenStudy (anonymous):

we now sub u by x 2+u-3-u/(2+u)^2 4x/1-x^2 =u -1=(2+(2+u)^2)

OpenStudy (anonymous):

no its wrong

OpenStudy (anonymous):

now we devertive this -1=(2+(2+u)^2) 0=3+(u^2+2u+4) -7=u^2+2u

OpenStudy (anonymous):

wrong